OpenStudy (anonymous):

Jane is traveling to Europe this summer with the French club. She plans to bring $300 to spend while she is there. If $90 in U.S. currency is equivalent to 63 euros, how many euros will she receive when she exchanges her money?

OpenStudy (calculusfunctions):

You can use ratios here. Do you know how?

OpenStudy (anonymous):

exchange rate is \[\frac{63}{90}=\frac{7}{10}=.7\]

OpenStudy (calculusfunctions):

@satellite73 you should know better than to give away the answer.

OpenStudy (anonymous):

i got 630 but i don't think i got it right

OpenStudy (anonymous):

did you multiply \(.7\times 300\)?

OpenStudy (anonymous):

no but i see now it should hav been something i did

OpenStudy (anonymous):

sorry about giving the answer, but since you asked i thought it might be what you were looking for

OpenStudy (anonymous):

thanks ok i needed that to understand